The market demand curve for a pair of duopolists is given as P=100- Q where Q= Q1+ Q2. The constant per unit marginal cost is 0 for firm 1 and c for firm 2 where c is some number. Find the equilibrium price, quantity and profit for each firm in the Bertrand model as a function of c

a. Equilibrium price equals P=0. Equilibrium quantity is Q1=Q2=10 with both earning Π12=0. Which one is correct?

---C= 0 OR C>0

b. Firm 2 produces zero Q2=0 and earns Π2= 0. Firm 1 charges P=c and produces Q1=100-c earning Π1=100 c – c2

Which one is correct?

---C= 0 OR C>0

Is C=0 or C>0 for the two questions
Answer #1

Under Bertrand competition, firms choose the price at which they want to sell their goods, and the market determines the quanFor firm 2, P=c c=100-9-22 Q2 =100-Q,-c Calculate the profits for the two firms as follows: Profits for Firm 1 = (P-MC)Q =(0-If equilibrium price equals 0, and quantities sold equal 10 each, and the profits of the two firms are $0, then the price hasThus, c= 0 is true.